A small open economy is described by the following equations:


C=50+0.75 (Y-T) I=200-20r NX=200-50e M/P=Y-40r

G=100 T=100 M=2,000 P=2 r*=4


(a) Derive and graph the IS* and LM* curves (Mundell-Fleming model)

(b) Calculate (i) the equilibrium rate and (ii) output

(c) Assume a floating exchange rate. Calculate what happens to (i) the exchange rate, (ii) income, (iii) net exports, and (iv) the money supply if the government raises taxes by 50. Use a graph to explain what you find.


1
Expert's answer
2021-05-24T13:01:05-0400

A small open economy is described the following equations.


"C = 50 + 0.75 (Y - T)
\\
T = 200
\\
I = 200 - 20r\\M/P = Y - 40r


\\
M = 3000
\\
P = 3
\\
r = 5
\\
G = 200
\\
NX = 200 - 50E"


a).

Equation for IS is given by Y = C + I + G + NX


"Y = 50 + 0.75 (Y - 200) + 200 - 20r + 200 + 200 - 50E
\\
Y = 2000 - 80r - 200E"


Equation for LM is


"Y - 40r = 3000
\\
Y = 40r + 1000"


(b) When "r = 0, Y = 2000-200E"


"0=2000-80r-200E


\\
80r=2000-2.5e"

"LM\Rightarrow y=3000+40r"


     When "r=0, y=3000"

     When "y=0, r=\dfrac{-3000}{40}=-75"


(c) At Equilibrium, IS=LM


  "2000-80r-200E=3000+40r"


   At r=5, 


  "2000-80(5)=3000+40(5)\\1400=200E\\E=7

\\[9pt]

 Y=2000-80(5)-200(7)=200"



(d) When G becomes 250, equation of IS becomes

Y = 550 + 0.75Y - 20r - 50E

Y = 2200 - 80r - 200E


LM equation does not change so we still have Y = 40r + 1000. Multiplier is 1/1-0.75 or 4 so income rises by 50*4 = $200. If Y = 1200 + 200 = 1400, we have

1400 = 40r + 1000, r = 10%. Hence at r = 10% and Y = 1400, we see that

1400 = 2200 - 80*10 - 200E

E* = 0.

At these levels, net exports = 200 - 50*0 = 200. There is no change in money supply